OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (mathstudent55):

Do you know how to start?

OpenStudy (anonymous):

For the first step we can imagine that we are multiplying the full expression (x+2) by 5, thus it is the same as saying (5times x) + (5times2)

OpenStudy (anonymous):

I would multiply the 5 through the parenthesis, and the 3 as well. You should now have 3-5x-10= 3x+3+6. Now combine like terms 3-10=-7, and 6+3=9. You should now have -5x-7=3x+9. Now bring constants to the right and variables with coefficients to the left(remember if its addition you need to subtract to cancel it out but you must also do the subtraction to the other side). You should have -8x=16. Now to get x by itself divide by -8 on both sides(division undoes multiplication). You should now have x=-2